OpenStudy (anonymous):

Solve the equation by completing the square. If nessecary, round to the nearest hundreth x^2-18x=19

OpenStudy (anonymous):

x^2-18x-19=0

OpenStudy (anonymous):

(x-9)^2-100

OpenStudy (anonymous):

\[(x-9)^{2}-100\]

OpenStudy (mertsj):

half of 18 is 9 and 9^2 is 81 so add 81 to both sides: \[x^2-18x+81=19+81\]

OpenStudy (mertsj):

Now factor the left side: \[(x-9)^2=100\]

OpenStudy (mertsj):

Now take the square root of both si: \[x-9=\pm10\]

OpenStudy (mertsj):

Now add 9 to both sides: \[x=10+9=19, x=-10+9=-1\]

OpenStudy (anonymous):

only 1;19 -1;19 3;6 and -3;1

OpenStudy (anonymous):

wait so its-1 19??

OpenStudy (mertsj):

The answers are x=19 x=-1

OpenStudy (anonymous):

thank you can you help me with one more?

OpenStudy (anonymous):

Which model is most appropriate for the data shown in the graph below?

OpenStudy (mertsj):

Looks like a parabola with vertex (-1,0). What do you think?

OpenStudy (anonymous):

i think its an exponential

OpenStudy (mertsj):

What are your choices?

OpenStudy (anonymous):

quadratic linear exponential and line

OpenStudy (mertsj):

If it is a parabola, its equation should be y = a(x-h)^2+k We know (h,k) is (-1,0) and we see that (1,1) is on the graph. So let's calculate the value of a and see if the other given points fit. If not, then it's an exponential.

OpenStudy (mertsj):